A coin rushes 10 times. What are the chances of 3 to 6 rounds? OA) 0,5524 OB) 0,4587 C) 0,7734 OD) 0,2289 O E) 0,7623
Answer #1


Let p be the probability success Then p=0.5 A coin rushes 10 times n= 10 q=1-p=1-0.5=0.5 Then the probability of getting x ch